Prove that the sequence {an/n} is convergent where { an} is a bounded sequence

1
Expert's answer
2022-01-25T12:24:08-0500

Due to the definition of the bounded sequence, there is "M>0" such that: "|a_n|<M" "\forall n\in{\mathbb{N}}". Therefore, "|\frac{a_n}{n}|<\frac{M}{n}". We receive: "|\frac{a_n}{n}|<\frac{M}{n}\rightarrow0". Thus, the sequence is convergent.